OpenStudy (love_to_love_you):

Please help me! Fan, medal, testimony! Urgent! Find the value of the variable. Show all work. [ x-5 9 ] = [ -7 w+1 ] [ 4 t+2 ] = [ 8-r 1 ]

OpenStudy (anonymous):

x-5 = 7 w+1 = 9 4 = 8 - r t +2 = 1 Find the values

OpenStudy (love_to_love_you):

@sangya21 can you help me with another?

OpenStudy (love_to_love_you):

@bohotness can you help?

OpenStudy (bohotness):

yes

OpenStudy (anonymous):

Sure

OpenStudy (love_to_love_you):

ok let me get it real quick

OpenStudy (bohotness):

ok

OpenStudy (love_to_love_you):

I need help on 13 and 15

OpenStudy (bohotness):

ifk abou t hat

OpenStudy (love_to_love_you):

ok, then can you help me sangya?

OpenStudy (anonymous):

15. [-14 -4] [33 -7] and 13 might take time

OpenStudy (love_to_love_you):

I already have the answers to them, the thing I need to do is show steps.

OpenStudy (love_to_love_you):

13. Undefined. 15. [-14 -2] [ 43 -7]

OpenStudy (anonymous):

13. Matrix A and B cannot be multiplied together because the number of columns in A does not equal the number of rows in B. In this case, the multiplication of these two matrices is not defined.

OpenStudy (anonymous):

15. AC - BD AC 1ST TERM OF AC = (6)(-2) + (1)(4) +(0)(2) + (8)(1) 2ND TERM OF AC = (6)(1) + (1)(0) +(0)(2) + (8)(1) 3RD TERM OF AC = (-4)(-2) + (3)(4) +(7)(2) + (11)(1) 4TH TERM OF AC = (-4)(-2) + (3)(4) +(7)(2) + (11)(1) BC 1ST TERM = (1)(5) + (3)(3) 2ND TERM = (1)(-2) +(3)(6) 3RD TERM =(-2)(5) + (4)(3) 4TH TERM =(-2)(-2) +(34(6) Do this in matrix form then subtract respective terms

OpenStudy (love_to_love_you):

ok, so..:

OpenStudy (love_to_love_you):

[(6)(-2) + (1)(4) +(0)(2) + (8)(1) (6)(1) + (1)(0) +(0)(2) + (8)(1)] [(-4)(-2) + (3)(4) +(7)(2) + (11)(1) (-4)(-2) + (3)(4) +(7)(2) + (11)(1)] MINUS [(1)(5) + (3)(3) (1)(-2) +(3)(6)] [(-2)(5) + (4)(3) (-2)(-2) +(34(6)]

OpenStudy (love_to_love_you):

=[-12 + 4 + 0 + 8 6 + 1 + 0 +8 ] [8 + 12 + 14 + 11 8 + 12 + 14 + 11] MINUS [5 + 9 -2 + 18] [-10 + 12 4 + 204] now what?

OpenStudy (love_to_love_you):

do I add? and then subtract the two?

OpenStudy (anonymous):

yups

OpenStudy (love_to_love_you):

=[0 15 ] [55 45] MINUS [14 16] [2 208]

OpenStudy (love_to_love_you):

is that right?

OpenStudy (anonymous):

sorry the last term of bd is (3)(6) thus making whole equation =[0 15 ] [55 45] MINUS [14 16] [2 22 sorry for late reply site prob